Leckwith Droves Allotments Association 

Minutes of Committee Meeting 

Sunday 2nd May 2021 at 12:00pm

Present: Nigel Bevan, Mitchell Young, Janis and Peter Sansom, Dorothy Hursthouse, Tony Tanti, Jimmy Pride, Paul Sullivan,  Gavin Michael.  

 

1. Apologies for absence:  Kim Stokes, Steve Stokes, Liz Irvine Rhys Livsey, Peter Edwards

 

2.Agree the minutes of the last meeting held on 28. 3. 2021:  Agreed by all present.

 

3. Matters arising.

  • Main gate lock has been replaced but discussions are ongoing with council to obtain an extra one for emergency use.

  • Veterans project: Discussions still ongoing regarding details of plot allocation. Peter S and JS liaising with council.

  • A complaints procedure has been set up on the website. 

  • Beekeeping project: Still in discussion with council over the possibility of finding an otherwise unlettable plot. Peter S and JS liaising with council.

 

4. Progress updates:

 

  1. Lettings report: Work is ongoing to prepare plots for letting. A total of 29 plots have been let since Oct 2020 and there are currently 6 vacant ones,5 of which need to have the soil improved before letting. Peter S requested that the council apply weedkiller and then the committee arrange for topsoil to be added. Purchase of this to be approximately £2,500. It was agreed that this money can come out of income from lettings. 53 plots referred to council for non-cultivation and negotiations are ongoing.

 

106 money: The committee discussed various projects that this could be spent on and agreed that the most important are: The pond renovation (£20,000); Purchase of 50 large water storage containers (£4,800); and an industrial chipper for supervised communal use (£2,800)

Action: MY to contact council to check these are an acceptable  use of 106 money

 

  1. Plot numbering. All the plots have now been numbered in a clear and logical manner to assist any deliveries and visitors to the site. Thanks to Peter S and JP. Celia Hart has been notified and would like committee to match these up to the numbers held by the council 

Action: JS and KS to organise this and draw up a map.

 

  1. Allotment Handbook. The text has been agreed by the committee and MY has agreed to format it for printing as a booklet. It was agreed that paper copies can be given to new plot holders and existing plot holders to be offered a electronic version.

Action: DH to send text to MY who will investigate formatting and printing and translating.

NB and  JS to look at compiling an email data base for plot holders. NB to compile a list of plot numbers for committee members to create a list for display near the shop.

 

  1. Pest control: third visit of this block due on 5th May

  2. Council will be visiting May 4th 9am with a view to arranging removal of asbestos. 

Action NB to be on site to meet council representative

 

  1. Communal area:Lime plaster has been purchased for Pizza oven. JP to complete the work.  PeterS and JP to replace membrane in communal area and cover with chippings. GM has tidied up the area beside the shop. It was agreed to purchase a replacement cover for the communal polytunnel ( approx£100) so it can be used by any plot holders who wish.

  2. Rubbish removal 4 loads of plastic and scrap metal has been removed from the site and paid for. 'Recycling area' can now be cleared and used for parking. Hopefully this will discourage dumping of rubbish. Any items for recycling can be advertised on the facebook group – this has worked successfully. Similarly it was decided not to create communal areas for manure and chippings delivery as this also attracts dumping.

  3. Social events: KS has suggested the possibility of organising a plant sale and social to test the pizza oven; a meal out for committee and friends of the committee; and a visit to the Malvern Show. All to be discussed when the regulations regarding covid 19 are more definite.

  4. Bio Diversity group. A plot holder is interested in starting such a group.

Action: KS to feedback that committee are happy for this to go ahead.

 

 

5. Any other business: 

  • Key difficulties on main gate: TT reported that one or two people are experiencing difficulty locking and unlocking new lock. He will investigate making the keys easier to use

  • Security cameras and signs: PS reported that the camera feed has been out of use temporarily. It has been suggested we put up further cameras to deter flytipping. 

Action:JS to investigate cost of professional signs to let people know cameras are around as well as signs warning against bringing of forbidden items to the site. Also signs for car parking areas.

 

Date and time of next meeting.

Our next meeting will be on Sunday 5th June  

 

Meeting finished at 1:30pm